(b) Find the impedance of a series RLC circuit if the inductive reactance, capacitive

Draw the corresponding voltage versus time graph. [10 marks]

reactance and resistance are 184 Ω, 144 Ω and -j30 Ω respectively. Also calculate the

(a) Write down the equation for a sinusoidal voltage of 50 Hz and its peak value is 20 V.

phase angle between voltage and current.


1
Expert's answer
2021-10-07T03:15:10-0400

(b) The impedance is


"Z=\sqrt{184^2+(144-30)^2}=216.5\Omega."

(a) The equation is


"v(t)=20\sin(100\pi t+\phi)"

The graph for phase shift of zero is below: